Prove the following identities:
sin^4 x - sin^2 x = cos^4 x - cos^2 x

Answer by Alan3354(69443)

Prove the following identities:
sin^4 x - sin^2 x = cos^4 x - cos^2 x
-------------
sin^2(x)*(sin^2(x) - 1) = cos^2(x)*(cos^2(x) - 1)
sin^2*(-cos^2) = cos^2*(-sin^2)
sin^2*cos^2 = sin^2*cos^2